Develop a spreadsheet to determine the next present value or present worth of the following project:
Investment: 170,000
Revenue/Savings: 32,000
Incremental Expense/Cost: 6,000
Salvage Value: 20,000
Project Life: 10 years
MACRS Schedule: 7 years
Tax Rate: 24%
MARR: 12%
Inflation: 3%
Complete this on the basis of a real dollars basis rather than actual dollars.
Is this a good investment to make?

Year Percent
1 14.29%
2 24.49%
3 17.49%
4 12.49%
5 8.93%
6 8.92%
7 8.93%
8 4.46%
Answer #1

Answer: -

According to question there are following data are given,

Investment: 170,000

Revenue/Savings: 32,000

Incremental Expense/Cost: 6,000

Salvage Value: 20,000

Project Life: 10 years

MACRS Schedule: 7 years

Tax Rate: 24%

MARR: 12%

Inflation: 3%


E-D A F=B-C-E G=F24% | Tax @ 24% Net Income Year Investment YO Y1 Y2 Revenue Expense MACRS Depreciation (170,000.00) 24,293.